# On $p$-adic Asai $L$-functions of Bianchi modular forms at non-ordinary primes and their decomposition into bounded $p$-adic $L$-functions

### v2

@@ 1. Introduction
This article addresses the next natural step: the construction of a $p$-adic Asai $L$-function when $\Psi$ is not ordinary at $p$. Our construction is also independent of the conjectural Asai motive over $\mathbb{Q}$. We attach a $p$-adic Asai $L$-function with unbounded denominators $L^{\mathrm{As}}_{p}(\Psi)$ to a $p$-non-ordinary small slope cuspidal Bianchi modular form $\Psi$ such that it interpolates critical values of the Asai $L$-function associated with $\Psi$. This construction generalizes the result of Loeffler–Williams [21, Theorem 7.5] from the $p$-ordinary case to the $p$-non-ordinary case.
@@
Roughly, Loeffler–Williams constructed a $p$-adic measure $L^{\mathrm{As}}_{p}(\Psi)$ using Asai–Eisenstein elements only for $j=0$. Since they work in the $p$-ordinary setting, by a deep theory of Kings implies that this measure is independent of the twist $j$. In particular, for a Dirichlet character $\theta$ of $p$-power conductor, by integrating $x^{j}\theta(x)$ against $L^{\mathrm{As}}_{p}(\Psi)$ recoveres the special $L$-value $L^{\mathrm{As}}(\Psi,\overline{\theta},j+1)$. See [21, Propositions 5.5, 6.3] for more details. We will briefly outline their work in Section 4.4. The method of using the only Asai–Eisenstein element for $j=0$ breaks down in $p$-non-ordinary setting.
@@
To circumvent this issue, we use the approach taken by Amice–Velu in [1] and Višik in [27]: interpolation of twists using polynomials. In this article, we obtain the $p$-adic Asai $L$-function $L^{\mathrm{As}}_{p}(\Psi)$ by constructing certain novel family of polynomials $P_{r,j}(T)\in E[T]$ of degree $p^{r-1}$, where $\Psi$ is a $p$-non-ordinary weight $(k,k)$ cuspidal Bianchi modular form, $r\in\mathbb{Z}_{\geq 1}$ and $0\leq j\leq k$. These polynomials are constructed by pairing an appropriate modular symbol associated to $\Psi$ with the Asai-Eisenstein elements constructed by Loeffler–Williams in [21, Sections 3, 4]. Furthermore, if we evaluate these polynomials at some special elements, we obtain the critical $L$-values $L^{\mathrm{As}}(\Psi,\overline{\theta},j+1)$ with some explicit factor. The polynomials $P_{r,j}$ satisfy certain norm and congruence properties. To prove the congruence properties of polynomials, we state and prove some new congruence results associated to Asai–Eisenstein elements. Then, by applying the techniques developed by Amice–Vélu, Višik, Perrin-Riou, and Büyükboduk–Lei to these polynomials, we construct the $p$-adic distribution with unbounded coefficients and some growth, that is, an element in the distribution space $\mathcal{H}_{E,v_{p}(a_{p})}(\Gamma)$, where $v_{p}(a_{p})$ is the slope of the ’$p$-th’ Fourier coefficient. We hope that one may use the $p$-adic distribution $L^{\mathrm{As}}_{p}(\Psi)$ to formulate and prove the Harron–Pottharst [11] style Iwasawa main conjecture over the distribution space, relating the ideal generated by this distribution over the distribution space to a certain Selmer complex. We briefly outline the construction of these polynomials and congruences in Subsection 1.2.
@@ 1.2. Outline of the construction of $P_{r,j}$
We also define a polynomial when $r=1$. See Equation (6.4). Here $Y^{*}_{F,1}(\mathfrak{N})$ is a locally symmetric space of a certain level (see 2 for the definition). Here $V_{k,k}$ are $T_{k,k}$ are some specific weight $(k,k)$-coefficient modules. The Asai-Eisenstein element $\prescript{}{c}{\Xi}^{k,j}_{p^{r},\mathfrak{N},at}$ is related to an Eisenstein series of weight 2k2j+22k-2j+2.These elements are ’the Betty counterparts’ of the known Euler system elements, such as Beilinson–Flach Euler systems. They satisfy certain norm-compatibilities like Beilinson–Flach elements. See [21, Theorem 3.13] for more details.
@@
For the definition of the moment map $\mathrm{mom}$, see Subsection 4.3.1. The Lemma 5.7 is the Betti cohomology analogue of [22, Theorem 3.3.5]. Furthermore, we prove another novel lemma :
@@ 1.3. Signed $p$-adic Asai $L$-functions
where $\tilde{Q}=\begin{pmatrix}\tilde{\alpha}&-\tilde{\beta}\\[5.0pt] -\alpha^{2}_{\overline{\mathfrak{p}}}p^{k-1}&-\alpha^{2}_{\overline{\mathfrak{p}}}p^{k-1}\end{pmatrix}$, and $\tilde{\underline{M}}$ satisfies properties from Proposition 7.2.
@@ 1.4. Plan of the article
In Section 2, we recall the definitions of locally symmetric spaces in which we are interested. We also define Hecke correspondences on these locally symmetric spaces. We define the Asai $L$-function of Bianchi modular forms in Section 3. Moreover, we recall the definitions of Bianchi modular forms and Bianchi modular symbols in this section. In Section 4, we describe the weight $2$ and higher weight Asai-Eisenstein elements constructed by Loeffler–Williams in [21]. The construction of these elements is closely related to the elements constructed in [17]. One can think of the Asai-Eisenstein elements as a Betti counterpart of known Euler system elements like Beilinson–Flach elements. We do some explicit calculations in Section 5 similar to [14, Theorem 6.2.4] and [17, Theorem 8.1.4]. We also prove some interpolation and congruence theorems, which are Betti analogues of the theorems in [22, Section 3.3]. In Section 6, we prove Theorem A using the interpolation of twists method of Amice–Velu, Perrin-Riou, and Büyükboduk–Lei. In the last section, we first recall the definition and construction of logarithmic matrices briefly. Then we prove Theorem B
@@ 3.1. Bianchi modular forms
We recall definitions of Bianchi modular forms and their Fourier expansions. For an integer $n\geq 0$, and a ring $R$, define $V_{n}(R)$ to be the space of homogeneous polynomials of degree $n$ in two variables $X,Y$ with coefficients in $R$. The space $V_{n}(R)$ can also be described as $\mathrm{Sym}^{n}(R^{2})$. The group of matrices $\mathrm{GL}_{2}(R)$ acts on $V_{n}(R)$ both from the right and left, and we denote the corresponding space by $V^{(r)}_{n}(R)$ or by $V^{(\ell)}_{n}(R)$, respectively. See [21, Section 2A] for the precise definitions.
@@
where the Kirillov function $W_{f}(-,\Psi)$ is a locally constant function on $(\mathbb{A}_{F}^{f})^{\times}$ with support contained in a compact subset of $\mathbb{A}_{F}^{f}$, $W_{\infty}\colon\mathbb{C}^{\times}\to V_{2k+2}(\mathbb{C})$ be the real analytic function defined in [28, 1.2.1(v)], and $e_{F}\colon\mathbb{A}_{F}/F\to\mathbb{C}^{\times}$ denote the unique continuous character such that its restriction to $F\otimes\mathbb{R}$ is $x_{\infty}\mapsto e^{2\pi i\mathrm{Tr}_{F/\mathbb{Q}}(x_{\infty})}$ (Also look at [21, Theorem 2.9]).
@@ 3.2. The Asai $L$-function of a Bianchi modular form
This Asai $L$-function $L^{\mathrm{As}}(\Psi,\theta,s)$ is absolutely convergent for $\mathrm{Re}(s)$ sufficiently large (one can take $\mathrm{Re}(s)>k+3$) and has meromorphic continuation for all $s\in\mathbb{C}$. See [21, Section 2E] for more details. For $s$ in the half-plane of convergence, $L^{\mathrm{As}}(\Psi,\theta,s)$ can be written as an Euler product
@@
See [10, Section 3] for more details.
@@
The Asai $L$-function appearing in Definition 3.1 is an "imprimitive" $L$-function. We can define a "primitive" Asai $L$-function using automorphic representations attached to Bianchi modular forms. Another way to define the Asai $L$-function of a Bianchi modular form $\Psi$ is the $L$-function attached to the tensor induction to $\mathbb{Q}$ of Galois representation associated with $\Psi$. See [21, Section 2F] and [10, Sections 3, 4].
@@
[21, Lemma 2.11] Let $\Psi$ be a normalized Bianchi eigenform of level $\mathfrak{n}$ coprime to $p$ and let $\Psi^{\prime}$ be a $p$-stabilization of $\Psi$ of level $p\mathfrak{n}$. Let $\alpha=c(p\mathcal{O}_{F},\Psi^{\prime})$. Then
@@ 3.3. Bianchi modular symbols
Theorem 3.4 (Eichler-Shimura-Harder, [21, Theorem 2.18]).
@@
Thus from Theorem 3.4, we have $\eta_{\Psi}\in\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathbb{C}))$, corresponding to an eigenform $\Psi\in S_{k,k}(U_{F,1}(\mathfrak{N}))$. Fix an eigenform $\Psi\in S_{k,k}(U_{F,1}(\mathfrak{N}))$, and let $F^{\prime}/\mathbb{Q}$ be a finite extension, large enough, containing $F$ and the Hecke eigenvalues of $\Psi$, $\mathfrak{P}$ be a prime of $F^{\prime}$ above $p$, and $\mathcal{O}_{F^{\prime},(\mathfrak{P})}$ be the valuation ring of $F^{\prime}$ at $\mathfrak{P}$. From [21, Section 2H], we can regard $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathcal{O}_{F^{\prime},(\mathfrak{P})}))$ as an $\mathcal{O}_{F^{\prime},(\mathfrak{P})}$-lattice in $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N},V_{kk}(F^{\prime}))$. Moreover, $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathcal{O}_{F^{\prime},(\mathfrak{P})}))$ is preserved by the action of Hecke operators $(T_{\mathfrak{a}})^{*}$ and $(U_{\mathfrak{a}})^{*}$.
@@
From [21, Proposition 2.20], we know that there exists a complex period $\Omega_{\Psi}\in\mathbb{C}^{\times}$, such that the quotient $\phi_{\Psi}\coloneqq\dfrac{\eta_{\Psi}}{\Omega_{\Psi}}$ forms an $\mathcal{O}_{F^{\prime},(\mathfrak{P})}$-basis of $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathcal{O}_{F^{\prime},(\mathfrak{P})}))[\Psi]$.
@@ 4. Asai-Eisenstein elements
In this section, we first recall definitions of modular units and Kato’s Siegel units. After that, we recall the definitions and constructions of weight $k=2$ as well as higher weight $k>2$ Asai-Eisenstein elements from [21, Sections 3, 5]. These are compatible families of classes in the Betti cohomology of locally symmetric spaces of Bianchi modular forms. Note that there is no étale cohomology in the Bianchi setting since the Bianchi manifolds, the locally symmetric spaces, are real manifolds of dimension $3$ and hence are not algebraic varieties. Loeffler–Williams constructed these elements by pushing forward Kato’s Siegel units to the Betti cohomology (under some maps) but using methods similar to the étale cohomology setting from [17].
@@ 4.1. Modular units and Siegel units
[12, Section 2.11, Proposition 3.11] The Siegel units are norm compatible, that is, if $N\mid N^{\prime}$ and $N$ and $N^{\prime}$ have the same prime divisors, then under the natural projection map $\mathrm{pr}:Y_{\mathbb{Q}}(M,N^{\prime})\to Y_{\mathbb{Q}}(M,N)$ we have
@@ 4.2. Weight $2$ Asai-Eisenstein elements
For a modular unit $u\in\mathcal{O}(Y_{\mathbb{Q},1}(N))^{\times}$, one can associate a Betti realization to $u$: $C(u)\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),\mathbb{Z})$. See [21, Proposition 3.2].
@@
    For more details, see the proof of [21, Theorem 3.13].

### v3

@@ 1. Introduction
This article addresses the next natural step: the construction of a $p$-adic Asai $L$-function when $\Psi$ is not ordinary at $p$. Our construction is also independent of the conjectural Asai motive over $\mathbb{Q}$. We attach a $p$-adic Asai $L$-function with unbounded denominators $L^{\mathrm{As}}_{p}(\Psi)$ to a $p$-non-ordinary small slope cuspidal Bianchi modular form $\Psi$ such that it interpolates critical values of the Asai $L$-function associated with $\Psi$. This construction generalizes the result of Loeffler–Williams [17, Theorem 7.5] from the $p$-ordinary case to the $p$-non-ordinary case.
@@
Roughly, Loeffler–Williams constructed a $p$-adic measure $L^{\mathrm{As}}_{p}(\Psi)$ using Asai–Eisenstein elements only for $j=0$. Since they work in the $p$-ordinary setting, by a deep theory of Kings implies that this measure is independent of the twist $j$. In particular, for a Dirichlet character $\theta$ of $p$-power conductor, by integrating $x^{j}\theta(x)$ against $L^{\mathrm{As}}_{p}(\Psi)$ recoveres the special $L$-value $L^{\mathrm{As}}(\Psi,\overline{\theta},j+1)$. See [17, Propositions 5.5, 6.3] for more details. We will briefly outline their work in Section 4.4. The method of using the only Asai–Eisenstein element for $j=0$ breaks down in $p$-non-ordinary setting.
@@
To circumvent this issue, we use the approach taken by Amice–Velu in [1] and Višik in [21]: interpolation of twists using polynomials. In this article, we obtain the $p$-adic Asai $L$-function $L^{\mathrm{As}}_{p}(\Psi)$ by constructing certain novel family of polynomials $P_{r,j}(T)\in E[T]$ of degree $p^{r-1}$, where $\Psi$ is a $p$-non-ordinary weight $(k,k)$ cuspidal Bianchi modular form, $r\in\mathbb{Z}_{\geq 1}$ and $0\leq j\leq k$. These polynomials are constructed by pairing an appropriate modular symbol associated to $\Psi$ with the Asai-Eisenstein elements constructed by Loeffler–Williams in [17, Sections 3, 4]. Furthermore, if we evaluate these polynomials at some special elements, we obtain the critical $L$-values $L^{\mathrm{As}}(\Psi,\overline{\theta},j+1)$ with some explicit factor. The polynomials $P_{r,j}$ satisfy certain norm and congruence properties. To prove the congruence properties of polynomials, we state and prove some new congruence results associated to Asai–Eisenstein elements. Then, by applying the techniques developed by Amice–Vélu, Višik, Perrin-Riou, and Büyükboduk–Lei to these polynomials, we construct the $p$-adic distribution with unbounded coefficients and some growth, that is, an element in the distribution space $\mathcal{H}_{E,v_{p}(a_{p})}(\Gamma)$, where $v_{p}(a_{p})$ is the slope of the ’$p$-th’ Fourier coefficient. We hope that one may use the $p$-adic distribution $L^{\mathrm{As}}_{p}(\Psi)$ to formulate and prove the Harron–Pottharst [10] style Iwasawa main conjecture over the distribution space, relating the ideal generated by this distribution over the distribution space to a certain Selmer complex. We briefly outline the construction of these polynomials and congruences in Subsection 1.2.
@@ 1.2. Outline of the construction of $P_{r,j}$
We also define a polynomial when $r=1$. See Equation (6.4). Here $Y^{*}_{F,1}(\mathfrak{N})$ is a locally symmetric space of a certain level (see 2 for the definition). Here $V_{k,k}$ are $T_{k,k}$ are some specific weight $(k,k)$-coefficient modules. The Asai-Eisenstein element $\prescript{}{c}{\Xi}^{k,j}_{p^{r},\mathfrak{N},at}$ is related to an Eisenstein series of weight 2k2j+22k-2j+2.These elements are ’the Betti counterparts’ of the known Euler system elements, such as Beilinson–Flach Euler systems. They satisfy certain norm-compatibilities like Beilinson–Flach elements. See [17, Theorem 3.13] for more details.
@@
For the definition of the moment map $\mathrm{mom}$, see Subsection 4.3.1. The Lemma 5.7 is the Betti cohomology analogue of [18, Theorem 3.3.5]. Furthermore, we prove another novel lemma :
@@ 1.3. Signed $p$-adic Asai $L$-functions
where $\tilde{Q}=\begin{pmatrix}\tilde{\alpha}&-\tilde{\beta}\\[5.0pt] -\alpha^{2}_{\overline{\mathfrak{p}}}p^{k+1}&-\alpha^{2}_{\overline{\mathfrak{p}}}p^{k+1}\end{pmatrix}$, and $\tilde{\underline{M}}$ satisfies properties from Proposition 7.2.
@@ 1.4. Plan of the article
In Section 2, we recall the definitions of locally symmetric spaces in which we are interested. We also define Hecke correspondences on these locally symmetric spaces. We define the Asai $L$-function of Bianchi modular forms in Section 3. Moreover, we recall the definitions of Bianchi modular forms and Bianchi modular symbols in this section. In Section 4, we describe the weight $2$ and higher weight Asai-Eisenstein elements constructed by Loeffler–Williams in [17]. The construction of these elements is closely related to the elements constructed in [16]. One can think of the Asai-Eisenstein elements as a Betti counterpart of known Euler system elements like Beilinson–Flach elements. We do some explicit calculations in Section 5 similar to [12, Theorem 6.2.4] and [16, Theorem 8.1.4]. We also prove some interpolation and congruence theorems, which are Betti analogues of the theorems in [18, Section 3.3]. In Section 6, we prove Theorem A using the interpolation of twists method of Amice–Velu, Perrin-Riou, and Büyükboduk–Lei. In the last section, we first recall the definition and construction of logarithmic matrices briefly. Then we prove Theorem B
@@ 3.1. Bianchi modular forms
We recall definitions of Bianchi modular forms and their Fourier expansions. For an integer $n\geq 0$, and a ring $R$, define $V_{n}(R)$ to be the space of homogeneous polynomials of degree $n$ in two variables $X,Y$ with coefficients in $R$. The space $V_{n}(R)$ can also be described as $\mathrm{Sym}^{n}(R^{2})$. The group of matrices $\mathrm{GL}_{2}(R)$ acts on $V_{n}(R)$ both from the right and left, and we denote the corresponding space by $V^{(r)}_{n}(R)$ or by $V^{(\ell)}_{n}(R)$, respectively. See [17, Section 2A] for the precise definitions.
@@
where the Kirillov function $W_{f}(-,\Psi)$ is a locally constant function on $(\mathbb{A}_{F}^{f})^{\times}$ with support contained in a compact subset of $\mathbb{A}_{F}^{f}$, $W_{\infty}\colon\mathbb{C}^{\times}\to V_{2k+2}(\mathbb{C})$ be the real analytic function defined in [22, 1.2.1(v)], and $e_{F}\colon\mathbb{A}_{F}/F\to\mathbb{C}^{\times}$ denote the unique continuous character such that its restriction to $F\otimes\mathbb{R}$ is $x_{\infty}\mapsto e^{2\pi i\mathrm{Tr}_{F/\mathbb{Q}}(x_{\infty})}$ (Also look at [17, Theorem 2.9]).
@@ 3.2. The Asai $L$-function of a Bianchi modular form
This Asai $L$-function $L^{\mathrm{As}}(\Psi,\theta,s)$ is absolutely convergent for $\mathrm{Re}(s)$ sufficiently large (one can take $\mathrm{Re}(s)>k+3$) and has meromorphic continuation for all $s\in\mathbb{C}$. See [17, Section 2E] for more details. For $s$ in the half-plane of convergence, $L^{\mathrm{As}}(\Psi,\theta,s)$ can be written as an Euler product
@@
See [9, Section 3] for more details.
@@
The Asai $L$-function appearing in Definition 3.1 is an "imprimitive" $L$-function. We can define a "primitive" Asai $L$-function using automorphic representations attached to Bianchi modular forms. Another way to define the Asai $L$-function of a Bianchi modular form $\Psi$ is the $L$-function attached to the tensor induction to $\mathbb{Q}$ of Galois representation associated with $\Psi$. See [17, Section 2F] and [9, Sections 3, 4].
@@
[17, Lemma 2.11] Let $\Psi$ be a normalized Bianchi eigenform of level $\mathfrak{n}$ coprime to $p$ and let $\Psi^{\prime}$ be a $p$-stabilization of $\Psi$ of level $p\mathfrak{n}$. Let $\alpha=c(p\mathcal{O}_{F},\Psi^{\prime})$. Then
@@ 3.3. Bianchi modular symbols
Theorem 3.4 (Eichler-Shimura-Harder, [17, Theorem 2.18]).
@@
Thus from Theorem 3.4, we have $\eta_{\Psi}\in\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathbb{C}))$, corresponding to an eigenform $\Psi\in S_{k,k}(U_{F,1}(\mathfrak{N}))$. Fix an eigenform $\Psi\in S_{k,k}(U_{F,1}(\mathfrak{N}))$, and let $F^{\prime}/\mathbb{Q}$ be a finite extension, large enough, containing $F$ and the Hecke eigenvalues of $\Psi$, $\mathfrak{P}$ be a prime of $F^{\prime}$ above $p$, and $\mathcal{O}_{F^{\prime},(\mathfrak{P})}$ be the valuation ring of $F^{\prime}$ at $\mathfrak{P}$. From [17, Section 2H], we can regard $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathcal{O}_{F^{\prime},(\mathfrak{P})}))$ as an $\mathcal{O}_{F^{\prime},(\mathfrak{P})}$-lattice in $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N},V_{kk}(F^{\prime}))$. Moreover, $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathcal{O}_{F^{\prime},(\mathfrak{P})}))$ is preserved by the action of Hecke operators $(T_{\mathfrak{a}})^{*}$ and $(U_{\mathfrak{a}})^{*}$.
@@
From [17, Proposition 2.20], we know that there exists a complex period $\Omega_{\Psi}\in\mathbb{C}^{\times}$, such that the quotient $\phi_{\Psi}\coloneqq\dfrac{\eta_{\Psi}}{\Omega_{\Psi}}$ forms an $\mathcal{O}_{F^{\prime},(\mathfrak{P})}$-basis of $\mathrm{H}^{1}_{\mathrm{c}}(Y_{F,1}(\mathfrak{N}),V_{kk}(\mathcal{O}_{F^{\prime},(\mathfrak{P})}))[\Psi]$.
@@ 4. Asai-Eisenstein elements
In this section, we first recall definitions of modular units and Kato’s Siegel units. After that, we recall the definitions and constructions of weight $k=2$ as well as higher weight $k>2$ Asai-Eisenstein elements from [17, Sections 3, 5]. These are compatible families of classes in the Betti cohomology of locally symmetric spaces of Bianchi modular forms. Note that there is no étale cohomology in the Bianchi setting since the Bianchi manifolds, the locally symmetric spaces, are real manifolds of dimension $3$ and hence are not algebraic varieties. Loeffler–Williams constructed these elements by pushing forward Kato’s Siegel units to the Betti cohomology (under some maps) but using methods similar to the étale cohomology setting from [16].
@@ 4.1. Modular units and Siegel units
[11, Section 2.11, Proposition 3.11] The Siegel units are norm compatible, that is, if $N\mid N^{\prime}$ and $N$ and $N^{\prime}$ have the same prime divisors, then under the natural projection map $\mathrm{pr}:Y_{\mathbb{Q}}(M,N^{\prime})\to Y_{\mathbb{Q}}(M,N)$ we have
@@ 4.2. Weight $2$ Asai-Eisenstein elements
For a modular unit $u\in\mathcal{O}(Y_{\mathbb{Q},1}(N))^{\times}$, one can associate a Betti realization to $u$: $C(u)\in\mathrm{H}^{1}(Y_{\mathbb{Q},1}(N),\mathbb{Z})$. See [17, Proposition 3.2].
@@
    For more details, see the proof of [17, Theorem 3.13].